The average age of the 40 members of a computer science camp is 17 years. There are 20 girls, 15 boys, and 5 adults. If the average age of the girls is 15 and the average age of the boys is 16, what is the average age of the adults?
(A) 26 (B) 27 (C) 28 (D) 29 (E) 30

Answer by mananth(16949)

20 girls avg age =15
15 boys avg age =16
Total 40 avg age = 17
Let the average age of adults be x. 5 adults
Total age of all students / total students = 17
5x+15*16+20*15 / 40 = 17
5x+240+300=40*17
5x+540=680
5x=680-540
5x=140
x= 140/5
x= 28
The average age of adults = 28